# FedAGHN: Personalized Federated Learning with Attentive Graph HyperNetworks

### v1

@@ 4 Experimental Settings
Data Heterogeneity: The data heterogeneity is considered through two Non-IID scenarios: the pathological setting and the practical setting.
Baselines and Evaluation: The proposed FedAGHN is compared against 15 baselines, including the local training method, 2 traditional federated learning methods, and 12 personalized federated learning methods. For each method, we save the best model that achieves the best performance on the local validation set during the training process, and evaluate it on the local test set of each client. The average of test accuracy across all clients is calculated.
Implementation Details: In the specific experimental implementation, we use the SGD optimizer and set the local batch size to 64. The number of local training epochs is set to 5, and the local learning rate for models is set to 0.01. We consider 20 clients participating with the join rate of 1.
Settings Pathological setting Practical setting (Dir(0.1))
[…]
5 Results and Discussions
  • RQ2: Is FedAGHN still effective across various degrees of statistical heterogeneity? (Section 5.2) (Table 2)
5.1 Main Experiment
To verify the superiority of the proposed FedAGHN against the 15 baseline methods, we fairly test them over the aforementioned datasets and two different Non-IID settings. The best performances and second best ones are bolded and underlined respectively in Table 1.
Overall, FedAGHN outperforms almost all baselines and achieves the state-of-the-art regarding all cases across the two Non-IID scenarios. This might be because that FedAGHN benefits from full consideration of personalization at the layer-wise level and adaptively tuned client collaborative relationships. The only case where FedAGHN doesn’t achieve the first place is in the pathological setting of CIFAR10, where it takes the second place, narrowly trailing behind FedPAC. This may be due to FedPAC makes it easier to achieve feature alignment when both the number of model parameters and the variety of data sample labels are small.
The traditional FL methods FedAvg and FedProx perform worse than local training. While the local adaptation methods FedAvg-FT and FedProx-FT, remain competitive among various PFL approaches.
For personalized aggregation-based methods, CFL performs competitively in the practical setting of TINY but performs worse in other settings. This might be because that CFL collaborates with a few clients whose updates of personalized local model parameters are most similar, ignoring the general information from other clients, making it less adaptable to various scenarios. H-FedAMP achieves high performance in the pathological setting of TINY and it outperforms FedAMP in most scenarios, possibly because H-FedAMP calculates the relative collaboration weights for each collaborating client, and it has a more flexible calculation of collaboration weights than FedAMP.
This study finds that finer-grained aggregation may result in improved performance. pFedLA performs well in most cases, possibly due to its fine-grained personalized aggregation. However, pFedLA is not adaptable to different scenarios, possibly because its weight generation relies entirely on the optimization effects of hypernetworks’ model parameters and clients’ embeddings. FedALA performs worse than pFedLA in most cases, since it integrates information from both local and global models at a finer granularity but overlooks the selection of collaborative relationships with other clients. FedAGHN outperforms all the above personalized aggregation-based methods, possibly because it can explicitly model collaborative relationships and adaptively tune collaboration weights while performing finer-grained aggregation.
To further explore the superior performance mentioned above, we also conducted an ablation study detailed in Appendix D to verify the effectiveness of its components.
5.2 Different Degrees of Heterogeneity
To investigate the performance of FedAGHN under various degrees of data heterogeneity, we manipulate the parameter $\beta$ in Dir($\beta$) to generate heterogeneous data with different distributions on the CIFAR10 dataset. The experimental results of FedAGHN and baseline methods under various degrees of data heterogeneity are presented in Table 2. FedAGHN outperforms all baselines across various degrees of data heterogeneity, which might be owing to the consideration of fine-grained aggregation and personalization at layer-wise.
6 Analysis of Collaboration in FedAGHN
  • RQ3: Can the collaboration graphs in FedAGHN effectively capture the collaborative relationships among clients? (Section E.2) (Figure 4, Table 3)
  • RQ4: Can FedAGHN meet the personalized requirements for layers and show differences among different layers? (Section 6.2.2) (Figure 5)
  • RQ5: Can FedAGHN dynamically tune collaborative relationships during FL process? (Section 6.3) (Table 4)
Settings Dir(0.1) Dir(0.2) Dir(0.5)
Local 85.00±0.42 80.47±0.21 71.41±0.12
FedAvg 60.72±1.13 60.04±1.52 65.06±0.11
FedAvg-FT 87.54±0.24 84.95±0.34 79.51±0.16
FedProx 60.58±0.84 60.02±1.28 65.05±0.08
FedProx-FT 87.65±0.34 84.69±0.33 79.60±0.33
Ditto 86.77±0.27 84.19±0.32 79.11±0.43
FedRep 87.27±0.09 84.22±0.31 77.47±0.34
pFedHN 83.60±0.42 80.83±0.34 73.18±0.41
FedCP 87.78±0.27 85.25±0.42 79.57±0.16
FedPAC 87.59±0.17 85.39±0.10 79.35±0.66
CFL 87.17±0.16 83.96±0.18 77.13±0.25
FedAMP 85.07±0.43 80.57±0.33 71.77±0.23
H-FedAMP 86.01±0.32 82.64±0.64 75.77±0.55
pFedLA 88.10±0.30 84.97±0.16 79.37±0.56
FedALA 87.63±0.18 84.89±0.21 79.64±0.36
FedAGHN 88.32±0.29 85.45±0.29 79.85±0.52
Table 2: The accuracy (%) on CIFAR10 for different heterogeneity.
6.1 Case Study Setting
In this section, we establish a case study setting to investigate the collaborative relationship learned by FedAGHN. The Non-IID scenario is introduced following Xu et al. (2023). Therefore, clients within each group follow the same data distribution, while clients across groups have different dominant classes. In this context, there are obvious differences between clients within the same group and clients in different groups, so we can intuitively estimate whether the collaborative relationships among clients are appropriately captured.
[…]
Local 73.96±0.10 40.25±0.25 28.56±0.40 14.99±0.04
CFL 79.66±0.21 49.55±0.34 38.59±0.46 25.93±0.29
H-FedAMP 78.23±0.49 46.11±0.67 27.09±0.81 26.74±0.22
FedAGHN 81.09±0.30 51.14±0.27 39.12±0.48 26.83±0.14
Table 3: The accuracy (%) of compared methods in case study
Table 3 has shown the performance of these methods in case study setting. FedAGHN performs best on these datasets in this setting, which could be explained by the appropriate collaboration graphs analyzed in Section 6.2 and 6.3.
6.2 Analysis of Collaboration Graphs
6.2.1 Comparison with other methods
To evaluate the efficacy of personalized aggregation methods in capturing collaborative relationships, the visualization of collaboration graphs in CFL, H-FedAMP and FedAGHN are shown in Figure 4-a, 4-b and 4-c respectively.
As illustrated in Figure 4, FedAGHN can effectively capture collaborative relationships among clients, with small self-weighting value. Overall, the collaboration weights of FedAGHN generally align with the similarity of clients’ data distributions in the case study setting. And FedAGHN has clear weight differences both intra-group and inter-group. For the clients within a same group, there are significant weight differences among clients, which implies that it assigns higher weights to more beneficial clients. Besides, it is shown in Figure 4-a that all clients within a group are assigned the same collaboration weights in CFL, while clients outside the group do not participate in collaboration. Furthermore, H-FedAMP introduces a fixed self-weighting value, while assigning slightly higher weights to clients within the same group and lower weights to others.
6.2.2 Layer-wise collaboration graphs
To further investigate the layer-wise personalization in FedAGHN, we visualize the collaboration graphs corresponding to different model layers on CIFAR100. Figure 5 reveals significant differences in the collaborative relationships captured by the shallow and deep layers of FedAGHN.
In the shallow layers (Figure 5-a, 5-b), the model tends to select more general information, with no significant differences between groups. FedAGHN introduces the tunable attentive mechanism to capture the fine-grained collaborative relationships among clients, which enables it to focus more on valuable information in the shallow layers, without excluding clients with different data distributions. In the deep layers (Figure 5-c, 5-d), the model tends to select more specific information and prefers collaboration among clients within the same group. Significant differences can be observed between the groups, especially in the last model layer.
6.3 Analysis of Dynamic Collaboration
To demonstrate that the proposed FedAGHN can dynamically tune the client collaboration during the training process, Table 4 presents the changes in collaboration weights during the training process. Overall, the weight differences among collaborating clients continue to change during the training process, possibly indicating that the adaptive tuning of collaborative relationships may enhance FedAGHN’s performance. The self-weighting value gradually decreases during the FL process, indicating that clients prefer to collaborate with others. FedAGHN’s adaptive tuning of collaborative relationships can better meet the collaboration requirements at different stages. The detailed analysis is presented in Appendix E.
Client First Prior Middle Final
$\text{FedAGHN}_{\text{all}}$ Self 0.0291 0.0192 0.0136 0.0126
$\text{FedAGHN}_{\text{all}}$ Similar 0.1436 0.0758 0.0674 0.0733
$\text{FedAGHN}_{\text{all}}$ Other 0.0338 0.0471 0.0490 0.0480
$\text{FedAGHN}_{\text{shallow}}$ Self 0.0291 0.0228 0.0175 0.0155
$\text{FedAGHN}_{\text{shallow}}$ Similar 0.1270 0.0703 0.0598 0.0586
$\text{FedAGHN}_{\text{shallow}}$ Other 0.0369 0.0479 0.0502 0.0505
$\text{FedAGHN}_{\text{deep}}$ Self 0.0291 0.0155 0.0097 0.0097
$\text{FedAGHN}_{\text{deep}}$ Similar 0.1601 0.0814 0.0750 0.0880
$\text{FedAGHN}_{\text{deep}}$ Other 0.0307 0.0463 0.0478 0.0454
7 Conclusion
This work proposed FedAGHN, a personalized aggregation-based method, to continuously capture appropriate collaborative relationships for fine-grained personalized aggregation. At each communication round, FedAGHN maintains a set of collaboration graphs for each client and each layer, which explicitly model the client-specific and layer-wise collaboration weights. The collaboration weights are obtained by the newly proposed tunable attentive mechanism. Finally, the personalized initial models are obtained by aggregating parameters over the learned collaboration graphs. Extensive experiments have demonstrated the superiority of FedAGHN. We also performed experiments to analyze how FedAGHN uniquely captures collaborative relationships across different model layers, various stages of the FL process, and self-participation weight. For future work, we suggest further understanding the evolution of collaborative relationships during the FL process, not only for FedAGHN but also for other PFL methods.
Appendix
Appendix A Details of Related Work
A.1 Various PFL Approaches
The vanilla PFL methods, such as FedAvg-FT and FedProx-FT, employ local training of the global model to obtain personalized models, which can be classified as local adaptation methods. Meta-learning-based methods Jiang et al. (2019); Fallah et al. (2020) and multi-task-learning-based methods Marfoq et al. (2021) apply the principles of meta-learning and multi-task learning to PFL, aiming to enhance the model adaptation.
Regularization-based methods, such as pFedMe T Dinh et al. (2020) and Ditto Li et al. (2021b), have new ideas for using global model in PFL. These methods learn personalized models for each client and employ L2 regularization to constrain the distance between the personalized models and the global model. Moreover, knowledge distillation is also a feasible way to absorb model information. FedPAC Xu et al. (2023) is one of knowledge distillation-based methods, introducing global feature centroids for explicit local-global feature representation alignment.
Parameter decoupling methods are commonly employed to achieve better personalized or shared models. FedRep Collins et al. (2021) first trains personalized classification layers and then trains a feature extractor. FedCP Zhang et al. (2023b) handles global and personalized information in the features separately using a global head and a personalized head, employing conditional policies to differentiate between global and personalized information.
Previous personalized aggregation-based methods mainly focus on certain aspects of collaboration among clients. Our work uses collaboration graphs to model the fine-grained collaborative relationships, and the collaboration weights on graphs can be adaptively optimized with clients’ objectives during FL process.
A.2 Graph Learning
The collaboration graph can be used to represent the relationship between clients in FL. However, natural graph structures do not exist in many real-world scenarios, which requires graph structure learning on the server side to establish collaboration graphs among clients Fu et al. (2022). A common approach is to construct collaboration graphs based on the model parameters of different clients. SFL Chen et al. (2022) constructs a similarity matrix using the similarity of personalized model parameters, then it forms a graph with trainable parameters as edge weights, and optimizes the graph with regularization during local training at client side; pFedGraph Ye et al. (2023) designs optimization goals on the server side and proceeds to optimize the collaboration graph by solving a quadratic program.
Previous approaches treat the model as a whole entity when constructing and utilizing the collaboration graph, disregarding the differences between different model layers. However, it is crucial to learn specific collaboration graphs for each model layer. Our work learns collaboration graphs for different model layers and the collaboration graphs can be optimized via tunable attentive mechanism during FL process.
A.3 Hypernetworks
By optimizing both the embedding vectors and hypernetwork parameters, pFedHN Shamsian et al. (2021) facilitates the generation of improved personalized model parameters. pFedLA Ma et al. (2022) validates that the aggregation parameters generated by hypernetworks can capture fine-grained relationships among clients, thereby enhancing the performance of personalized models. Additionally, hypernetworks are on the server side, which mitigates the introduction of additional communication overhead.
The hypernetwork architecture in pFedLA employs SoftMax layers to generate the weights with differences for different model layers. The majority of hypernetwork parameters are shared by all clients in pFedHN. The optimization process involves a larger number of parameters, which may potentially result in smaller differences between clients or between layers.
Regarding our work, we reduce the number of trainable parameters in the hypernetwork and retain only the vital training parameters which can tune the collaboration graphs, reducing the optimization burden while increasing the differences among model layers.
Appendix B Details of the Proposed FedAGHN
B.1 Pseudocode of the FedAGHN Algorithm
The pseudocode of the proposed FedAGHN algorithm is presented in Algorithm 1.
@@ B.1 Pseudocode of the FedAGHN Algorithm
B.2 Calculate Attentive Weights
@@
Note that, the two trainable parameters $q^{r}_{i}$ and $p^{r}_{i}$ are vital for $\text{AGHN}_{i}$. An increased value of $q_{i}^{r}$ denotes that client $i$ prefers greater difference among collaboration weights at layer $r$, while an increased value of $p_{i}^{r}$ strengthen the self-weight for client $i$ and thus reduce the inclination to cooperate with other clients.
B.3 Aggregate Parameters over Graphs
@@
It might be worth noting that attentive weights are usually calculated using node features $h$ when aggregating $h$ in traditional graph learning, which is different from utilizing another node features $a$ in our above implementation. In fact, this is exactly the unique change we make when applying the graph learning paradigm to federated learning. This is because the update of model parameters represents the updated direction of the model on local data in PFL, it can better reflect the similarity of data distributions among different clients than the model parameters, thereby better modeling collaborative relationships among clients.
B.4 The Update of AGHN
@@
The collaboration weights among clients should adapt to the changes over communication rounds in federated learning. For brevity, considering each of $R$ layers at client $i$, the two sets of trainable parameters are denoted as $q_{i}=\{q_{i}^{1},q_{i}^{2},...,q_{i}^{R}\}$ and $p_{i}=\{p_{i}^{1},p_{i}^{2},...,p_{i}^{R}\}$, where each layer at client $i$ maintains a pair of scalars $p_{i}^{r}$ and $q_{i}^{r}$.
@@
Following the previous works Shamsian et al. (2021); Ma et al. (2022), we accordingly adopt a more general way to update $p_{i}$ and $q_{i}$ based on gradient descent, their updating rules are as follows
@@
where $\eta_{hn}$ is the learning rate for hyper-networks; $\Delta{q_{i}}$ denotes the gradient vector with $R$ dimensions for client $i$’s updating; $\nabla_{q_{i}}$ is partial derivative with respect to $q_{i}$; the operator T indicates for transpose; $\bar{\theta}_{i}$ represents the personalized initial model parameters after aggregation; $\Delta\theta_{i}$ denotes the update of personalized local model parameters. It might be worth mentioning that there is no additional communication cost while calculating the trainable parameters $p_{i}$ and $q_{i}$.
Appendix C Details of Experimental Settings
C.1 Data Heterogeneity

### v2

@@ 3.2 FedAGHN Algorithm
3.3 Attentive Graph Hyper-Networks
In this section, we take the attentive graph hypernetwork for client $i$ (i.e., $\text{AGHN}_{i}$) as an example to elaborate the technical details. For simplicity, we now focus on one communication round and remove the superscript reflecting the round. The AGHN and its workings are illustrated in Figure 3. The internal computation of AGHN in a detailed pseudocode format is also shown in Algorithm 2.
Algorithm 2 Attentive Graph Hyper-Networks
Input: The model parameters $\Theta^{(t)}$ and the update of model parameters $\Delta\Theta^{(t)}$ in the round $t$.
Parameter: ${\text{AGHN}}_{i}$’s trainable parameters $p_{i}$ and $q_{i}$.
Output: The personalized initial model $\bar{\theta}^{(t+1)}_{i}$ of client $i$ for the next round $t+1$.
1:procedure ${\text{AGHN}}_{i}$($\Theta^{(t)},\Delta\Theta^{(t)};p_{i};q_{i}$)
2:  Set $h=\Theta^{(t)}$ and $a=\Delta\Theta^{(t)}$.
3:  for each layer $r\in\{1,\ldots,R\}$ do
4:   Initialize the collaboration graph $G_{i}^{r}$ ’s topology.
5:   Initialize of node features $(h^{r};a^{r})$:
6:   $h^{r}=\{h_{1}^{r},h_{2}^{r},...,h_{N}^{r}\}=\{\theta_{1}^{r},\theta_{2}^{r},...,\theta_{N}^{r}\}$
7:   $a^{r}=\{a_{1}^{r},a_{2}^{r},...,a_{N}^{r}\}=\{\Delta\theta_{1}^{r},\Delta\theta_{2}^{r},...,\Delta\theta_{N}^{r}\}$.
8:   Calculate the collaboration weights:
9:   $\widetilde{\alpha}_{ij}^{r}=\left\{\begin{matrix}\frac{{e^{q_{i}^{r}\text{cos}(a_{i}^{r},a_{j}^{r})}}}{{\textstyle\sum_{j}e^{q_{i}^{r}\text{cos}(a_{i}^{r},a_{j}^{r})}}}&\quad j\neq i\\ {p_{i}^{r}}&\quad j=i\end{matrix}\right$..
10:   Normalize the attentive weights:
11:   $\alpha_{ij}^{r}=\frac{\widetilde{\alpha}_{ij}^{r}}{{\textstyle\sum_{j=1}^{n}}\widetilde{\alpha}_{ij}^{r}}\quad\forall j\in\{1,2,\dots,N\}.$
12:   Aggregate parameters over graphs:
13:   ${\bar{\theta}}_{i}^{r}=\sum_{j\in\{1,2,\dots,N\}}\alpha_{ij}^{r}{h_{j}^{r}}.$
14:  end for
15:  Collect $\bar{\theta}^{(t+1)}_{i}=\{{\bar{\theta}}_{i}^{1},{\bar{\theta}}_{i}^{2},\dots,{\bar{\theta}}_{i}^{R}\}$.
16:  return $\bar{\theta}^{(t+1)}_{i}$
17:end procedure
The key principle of AGHNs is to learn the collaboration graphs (i.e., collaboration weights) such that each client can aggregate its beneficial model parameters from all clients based on the learned collaboration graphs. Because the layers of deep learning models may have different functionalities, we also consider different collaboration patterns for different layers respectively, such that $\text{AGHN}_{i}$ for client $i$ maintains $R$ collaboration graphs $G_{i}=\{G^{1}_{i},G^{2}_{i},\dots,G^{R}_{i}\}$ where $R$ is the number of layers in the model at client $i$. It should be noted that there is no collaboration among the different layers.
For convenience, we follow the graph learning terminology in describing the AGHN at client $i$ and layer $r$, which corresponds to the collaboration graph $G_{i}^{r}$.
3.3.1 Initialize collaboration graphs
For the initialization of the graph topology, the collaboration graph $G^{r}_{i}$ is established for the central node (client) $i$, such that there are totally $N$ nodes with edges connected to node $i$ from all nodes $j$ (including self-connection from $i$).
For the initialization of node features $(h^{r};a^{r})$, we take the personalized local model parameters $\Theta^{r}$ from clients at layer $r$ to obtain the set of node features $h^{r}$ as $h^{r}=\{h_{1}^{r},h_{2}^{r},\dots,h_{N}^{r}\}=\{\theta_{1}^{r},\theta_{2}^{r},\dots,\theta_{N}^{r}\}$ and we take the updates of personalized local model parameters $\Delta\Theta^{r}$ from clients at layer $r$ to obtain the set of node features $a^{r}$ as $a^{r}=\{a_{1}^{r},a_{2}^{r},\dots,a_{N}^{r}\}=\{\Delta\theta_{1}^{r},\Delta\theta_{2}^{r},\dots,\Delta\theta_{N}^{r}\}$ where the set of node features $h^{r}$ is employed for aggregation over graphs, whereas the set of node features $a^{r}$ is utilized to calculate attentive weights. The initial node features for the first round are randomly parameterized, whereas the initial node features for each of the other rounds come from the immediately preceding round.
3.3.2 Calculate attentive weights
Specifically, we calculate the collaboration weights between node $i$ and each node $j\in\{1,2,\dots,N\}$ via a newly proposed tunable attentive mechanism:
@@ 3.3.2 Calculate attentive weights
To optimize the attentive weights with respect to the client $i$’s objective, we introduce two trainable parameters to further tune the attentive weights. Concretely, $q^{r}_{i}$ is adopted to tune the attentive weights between central node $i$ and the other nodes, while $p^{r}_{i}$ is used to tune the attentive weight of self-connection of central node $i$. Note that, the two trainable parameters $q^{r}_{i}$ and $p^{r}_{i}$ are vital for $\text{AGHN}_{i}$-1-1-1Remarks of trainable parameters: An increased value of $q_{i}^{r}$ denotes that client $i$ prefers greater difference among collaboration weights at layer $r$, while an increased value of $p_{i}^{r}$ strengthen the self-weight for client $i$ and thus reduce the inclination to cooperate with other clients. The detailed analysis of $p$ and $q$ can be found in the experiments such as Section 5.2 and 6., because they can explicitly capture and tune the collaborative relationships according to the client’s objective in the end-to-end paradigm.
To accelerate convergence and avoid exploding gradients, the normalized attentive weights $\alpha_{i}^{r}=\{\alpha_{i1}^{r},\alpha_{i2}^{r},\dots,\alpha_{iN}^{r}\}$ for collaboration graph $G_{i}^{r}$ can be calculated via
$\alpha_{ij}^{r}=\frac{\widetilde{\alpha}_{ij}^{r}}{{\textstyle\sum_{j=1}^{n}}\widetilde{\alpha}_{ij}^{r}}\quad\forall j\in\{1,2,\dots,N\}.$ (8)
3.3.3 Aggregate parameters over graphs
The collaboration graph $G_{i}^{r}$ includes the attentive weights $\{\alpha_{i1}^{r},\alpha_{i2}^{r},\dots,\alpha_{iN}^{r}\}$ and the node features $h^{r}$ from the previous round. The personalized initial model $\bar{\theta}_{i}^{r}$ for client $i$ at layer $r$ can be obtained by weighted aggregating over $G_{i}^{r}$ as follows:
${\bar{\theta}}_{i}^{r}=\sum_{j\in\{1,2,\dots,N\}}\alpha_{ij}^{r}{h_{j}^{r}}.$ (9)
@@ 3.3.3 Aggregate parameters over graphs
It might be worth noting that attentive weights are usually calculated using node features $h$ when aggregating $h$ in traditional graph learning, which is different from utilizing another node features $a$ in our implementation described herein. In fact, this is exactly the unique change we make when applying the graph learning paradigm to federated learning. This is because the update of model parameters represents the updated direction of the model on local data in FL, it can better reflect the similarity of data distributions among different clients than the model parameters can, thereby better modeling collaborative relationships among the clients.
3.3.4 The update of AGHN
@@
The collaboration weights among clients should adapt to the changes over communication rounds in FL. For brevity, considering each of the $R$ layers at client $i$, the two sets of trainable parameters are denoted as $q_{i}=\{q_{i}^{1},q_{i}^{2},\dots,q_{i}^{R}\}$ and $p_{i}=\{p_{i}^{1},p_{i}^{2},\dots,p_{i}^{R}\}$, where each layer at client $i$ maintains a pair of scalars $p_{i}^{r}$ and $q_{i}^{r}$.
@@
Following earlier research studies Shamsian et al. (2021); Ma et al. (2022), we adopt a more general way to update $p_{i}$ and $q_{i}$ based on gradient descent, their updating rules are as follows
@@
where $\eta_{hn}$ is the learning rate for the hyper-networks; $\Delta{q_{i}}$ denotes the gradient vector with $R$ dimensions for client $i$’s updating; $\nabla_{q_{i}}$ is the partial derivative with respect to $q_{i}$; the operator T indicates the transpose operation; $\bar{\theta}_{i}$ represents the personalized initial model parameters after aggregation; and $\Delta\theta_{i}$ denotes the update of the personalized local model parameters. It might be worth mentioning that there is no additional communication cost while calculating the trainable parameters $p_{i}$ and $q_{i}$.
4 Experimental Settings
This section introduces the experimental settings, including datasets, models, data heterogeneity, baselines, evaluation, and implementation details. The more specific summary of settings is listed in Table 2 for better reproducibility.
Datasets & Models Details
CIFAR-10 4-layer CNN + BN, 10 classes
CIFAR-100 4-layer CNN + BN, 100 classes
CIFAR-100* ResNet-18, 100 classes
Tiny-ImageNet ResNet-18, 200 classes
Data Heterogeneity Details
Pathological Randomly assign 5/20/40 classes
Practical Dirichlet distribution $\beta=0.1$
Training Settings Details
Optimizer SGD
Local Batch Size 64
Local Epochs 5
Learning Rate 0.01 (models), 0.005 (hypernetwork)
Client Number 20
Comm. Rounds 200/80/40 (empirical convergence)
Split Ratio Train:Val:Test = 7:1:2 per client
Table 2: Comprehensive Experimental Settings
4.1 Datasets and Models
To evaluate the performance of FedAGHN, we perform a number of tests on three benckmark datasets: CIFAR10, CIFAR100 Krizhevsky et al. (2009), and Tiny-ImageNet Chrabaszcz et al. (2017). For CIFAR10 and CIFAR100, we use the 4-layer CNN architecture with Batch Normalization (BN) layers similar to that done in past research studies Ma et al. (2022). To assess the performance on larger backbone networks and datasets, we utilize ResNet-18 as the backbone model for CIFAR100 and Tiny-Imagenet. In the following sections, $\text{CIFAR100}^{*}$ denotes using ResNet-18 on CIFAR100, and TINY indicates using ResNet-18 on Tiny-ImageNet.
4.2 Data Heterogeneity
The data heterogeneity is considered through two Non-IID scenarios: the pathological setting and the practical setting.